# Floquet Engineering Magnetism and Superconductivity in the Square-Lattice Hubbard Model

### v1

@@ Abstract
We study the interplay of magnetic order and superconductivity in the square-lattice Hubbard model under periodic driving with circularly polarized light. Formulating diagrammatic techniques based on the random-phase approximation in terms of Floquet Green’s functions, allows us to analyze fluctuation-driven unconventional pairing for weak-to-moderate interactions. The interplay of repulsive interactions and photo-assisted hopping of electrons gives rise to a rich magnetic phase diagram featuring an antiferromagnetic-to-ferromagnetic crossover prior to a Floquet Lifshitz transition. Close to the antiferromagnetic transition, topological $d+id$-wave superconductivity prevails in the phase diagram for a wide range of drive parameters. At intermediate-to-high frequency driving near the Floquet Lifshitz transition, superconducting orders are tuned from spin-singlet $d$-wave to spin-triplet $p$-wave character, providing an effective protocol for Floquet engineering topological superconductivity.
@@
where ${\hat{c}}^{\dagger}_{i\sigma}$ ($\hat{c}_{i\sigma}$) is the creation (annihilation) operator of a fermion with spin $\sigma$ on site $i$ and $\hat{n}_{i\sigma}={\hat{c}}^{\dagger}_{i\sigma}\hat{c}_{i\sigma}$. $U$ and $\tilde{t}_{ij}$ are the on-site repulsion and hopping between site $i$ and $j$, respectively. For the latter, we only allow for nearest-neighbor hopping $\tilde{t}$ and next-nearest-neighbor hopping $\tilde{t}^{\prime}$. Circular polarized field, breaking time-reversal symmetry, is included via Peierls substitution [41] with its vector potential
@@
In the wide-band limit [42], this introduces a hybridization $\gamma=\pi\rho_{0}|V|^{2}$ with $\rho_{0}$ being the reservoir density of states.
We assume the system to be initially metallic and investigate how driving affects collective fluctuations and the emergence of magnetic and superconducting order in the weak-to-moderately interacting regime. Due to the weak coupling with the bath, we expect heating effects to be moderate and a Floquet steady state that is synchronized with the drive to be achievable (on the timescale $t^{*}\sim\gamma^{-1}$).
High-frequency limit—
@@
Floquet random phase approximation—
@@
[…]
@@
introducing the multi-index $K=(\omega,\mathbf{k})$ for the real frequency $\omega$ and the momentum $\mathbf{k}$. The Keldysh Green’s functions that fulfill $G_{\mathbf{k}}^{X}(t,t^{\prime})=G_{\mathbf{k}}^{X}(t+T,t^{\prime}+T)$ for $X\in\{\mathrm{R},\mathrm{A},<\}$ were transformed into
@@
with $\overline{t}=\frac{t+t^{\prime}}{2}$ and $\tau=t-t^{\prime}$ (see Supplementary Material (SM) for details [53]). Eq. (13) is a matrix in Floquet space $\hat{\chi}^{\mathrm{ph}}_{Q}$ and we can define the spin (s) and charge (c) susceptibility as a geometric sum within the random phase approximation
@@
Floquet induced response transition—
@@
Floquet engineering magnetic order—
@@
Superconductivity near the magnetic instability—
After establishing the modified magnetic phase diagram of the Floquet-driven Hubbard model, we continue to investigate spin-fluctuation mediated superconductivity. In Figs. 2(b)-(c), one can find different superconducting regimes in the vicinity of magnetic order for the low- and intermediate-frequency case. In the absence of driving ($E_{0}/\Omega=0$), but with coupling to the reservoir, we find spin-fluctuation induced $d_{x^{2}-y^{2}}$ pairing, as expected from equilibrium [60, 61, 62]. Upon turning on the drive, the $d_{x^{2}-y^{2}}$ ordering persists in the vicinity of Néel-type AFM ordering ($\mathrm{AFM}1$) for both low ($\Omega=2.0$) and intermediate ($\Omega=5.0$) frequencies up to $E_{0}/\Omega\lesssim 1$. Although the dominant pairing symmetry remains $d_{x^{2}-y^{2}}$, driving induces a mixing with $d_{xy}$ symmetry components (see inset). To investigate the mixing of the sub-components and their topological properties, we calculate the Berry curvature of the associated Bogoliubov-de-Gennes Hamiltonian using Fukui’s method [63]. We find that a topological $d_{x^{2}-y^{2}}+id_{xy}$ order with Chern number $C=2$ is induced by driving in the vicinity of AFM1 for both $\Omega=2.0,5.0$. In the low-frequency case (Fig. 2(b)), chiral $d$-wave superconductivity is suppressed as soon as the magnetic instability switches to FM order. Although triplet superconductivity could be expected to emerge in the vicinity of the FM instability, we do not find this order for the given parameters. However, in the intermediate frequency regime (Fig. 2(c)), triplet $p+ip$-wave superconductivity can be stabilized around $E_{0}/\Omega\sim 2$ close to the Floquet Lifshitz transition (see also the inset in Fig. 2(c)). With a further increase of $E_{0}/\Omega$, the effect of the dynamical localization point becomes evident, where $d_{x^{2}-y^{2}}+id_{xy}$ order is stabilized again, but now with the $d_{xy}$ component being dominant ($d\mathrm{SC}2$). This phase appears in the vicinity of the magnetic instability towards the stripe AFM order ($\mathrm{AFM}2$), rather than $\mathrm{AFM}1$. All orders are found to be topologically nontrivial (see SM [53]).
Discussion—
@@ Appendix F Floquet amplitude expansion
We evaluate Equation˜74 for $\omega=0$ and $\mathbf{q}=\mathbf{q}_{C}$, where $\mathbf{q}_{C}$ is determined from the exact calculation described in the main text. According to our analysis, $\mathbf{q}_{C}$ can also be accurately obtained from Equation˜74 up to $E_{0}/\Omega\leq 0.4$ for $\Omega=2$, $E_{0}/\Omega\leq 0.8$ for $\Omega=4$, $E_{0}/\Omega\leq 2$ for $\Omega=6$, and $E_{0}/\Omega\leq 2.2$ for $\Omega=8$.

### v2

@@ Abstract
We study the interplay of magnetic order and superconductivity in the square-lattice Hubbard model under periodic driving with circularly polarized light. Formulating diagrammatic techniques based on the random-phase approximation in terms of Floquet Green’s functions, allows us to analyze fluctuation-driven unconventional pairing for weak-to-moderate interactions. The interplay of repulsive interactions and photo-assisted hopping of electrons gives rise to a rich magnetic phase diagram featuring an antiferromagnetic-to-ferromagnetic transition prior to a Floquet Lifshitz transition. Close to the antiferromagnetic transition, topological $d+id$-wave superconductivity prevails in the phase diagram for a wide range of drive parameters. At intermediate-to-high frequency driving near the Floquet Lifshitz transition, superconducting orders are tuned from spin-singlet $d$-wave to spin-triplet $p$-wave character, providing an effective protocol for Floquet engineering topological superconductivity.
@@
where ${\hat{c}}^{\dagger}_{i\sigma}$ ($\hat{c}_{i\sigma}$) is the creation (annihilation) operator of a fermion with spin $\sigma$ on site $i$ and $\hat{n}_{i\sigma}={\hat{c}}^{\dagger}_{i\sigma}\hat{c}_{i\sigma}$. $U$ and $\tilde{t}_{ij}$ are the on-site repulsion and hopping between site $i$ and $j$, respectively. For the latter, we allow for nearest-neighbor hopping $\tilde{t}$ as well as next-nearest-neighbor hopping $\tilde{t}^{\prime}$ to break the perfect nesting condition and allow for superconducting order close to the magnetic transition [41]. Circularly polarized light, breaking time-reversal symmetry, is included via Peierls substitution [42] with its vector potential
@@
In the wide-band limit [43], this is modeled as a constant hybridization $\gamma=\pi\rho_{0}|V|^{2}$ with $\rho_{0}$ being the reservoir density of states. The (thermal) reservoir is defined by the reservoir distribution function $f(\omega)=(e^{(\omega-\mu)/T_{\mathrm{res}}}+1)^{-1}$ at temperature $T_{\mathrm{res}}$ and chemical potential $\mu$. The hybridization enters as an imaginary selfenergy through the Dyson equation into the reservoir dressed Green’s function of the system.
The system is initially in the uncorrelated normal state (paramagnetic), and we investigate how driving affects collective fluctuations and the emergence of magnetic and superconducting order in the weak-to-moderately interacting regime. The weak coupling to the bath facilitates the formation of a Floquet steady-state that is synchronized with the drive (on the timescale $t^{*}\sim\gamma^{-1}$[44]. In this formalism, dissipation and heating balance each other and no other heating channels are available.
High-frequency limit.
@@
Floquet random phase approximation.
@@
$\chi^{\mathrm{ph}}_{nn^{\prime}}(Q)=\frac{i}{N}\sum_{l,\mathbf{k}}\int\frac{\mathrm{d}\omega}{2\pi}\;$ $G^{\mathrm{R}}_{n,n^{\prime}+l}(K)G^{<}_{l,0}(K-Q)$
@@
introducing the multi-index $K=(\omega,\mathbf{k})$ for the real frequency $\omega$ and the momentum $\mathbf{k}$ as well as discrete Floquet indices $n,n^{\prime},l$. The Keldysh Green’s functions that fulfill $G_{\mathbf{k}}^{X}(t,t^{\prime})=G_{\mathbf{k}}^{X}(t+T,t^{\prime}+T)$ for $X\in\{\mathrm{R},\mathrm{A},<\}$ were Fourier transformed twice into
@@
with $\overline{t}=\frac{t+t^{\prime}}{2}$ and $\tau=t-t^{\prime}$ (see Supplemental Material (SM) for details [55]). Eq. (33) is a matrix in Floquet space $\hat{\chi}^{\mathrm{ph}}_{Q}$ and we can define the spin (s) and charge (c) susceptibility as a geometric sum within the random phase approximation
@@
Floquet induced response transition.
@@
Floquet engineering magnetic order.
@@
Superconductivity near the magnetic instability.
After establishing the modified magnetic phase diagram of the Floquet-driven Hubbard model, we continue to investigate spin-fluctuation mediated superconductivity. The possible symmetries of the superconducting order parameter are given by the irreducible representations of the $C_{4v}$ point group of the underlying square lattice. In addition to the $s$-wave, the irreps are given in momentum space by $\Delta_{d_{x^{2}-y^{2}}}(\mathbf{k})=\cos k_{x}-\cos k_{y}$, $\Delta_{d_{xy}}(\mathbf{k})=\sin k_{x}\sin k_{y}$ and $\Delta_{p_{x(y)}}(\mathbf{k})=\sin k_{x(y)}$. By projecting the superconducting gap onto the irreps we can determine the gap symmetry [41]. In Figs. 2(b)-(c), one can find different superconducting regimes in the vicinity of magnetic order for the low- and intermediate-frequency case. In the absence of driving ($E_{0}/\Omega=0$), but with coupling to the reservoir, we find spin-fluctuation induced $d_{x^{2}-y^{2}}$ pairing, as expected from equilibrium [62, 63, 64]. Upon turning on the drive, the $d_{x^{2}-y^{2}}$ ordering persists in the vicinity of Néel-type AFM ordering ($\mathrm{AFM}1$) for both low ($\Omega=2.0$) and intermediate ($\Omega=5.0$) frequencies up to $E_{0}/\Omega\lesssim 1$. Although the dominant pairing symmetry remains $d_{x^{2}-y^{2}}$, driving induces a mixing with $d_{xy}$ symmetry components (see inset). To investigate the mixing of the sub-components and their topological properties, we calculate the Berry curvature of the associated Bogoliubov-de-Gennes Hamiltonian using Fukui’s method [65]. We find that a topological $d_{x^{2}-y^{2}}+id_{xy}$ order with Chern number $C=2$ is induced by driving in the vicinity of AFM1 for both $\Omega=2.0,5.0$. In the low-frequency case (Fig. 2(b)), chiral $d$-wave superconductivity is suppressed as soon as the magnetic instability switches to FM order. Although triplet superconductivity could be expected to emerge in the vicinity of the FM instability, we do not find this order for the given parameters. However, in the intermediate frequency regime (Fig. 2(c)), triplet $p+ip$-wave superconductivity can be stabilized around $E_{0}/\Omega\sim 2$ close to the Floquet Lifshitz transition (see also the inset in Fig. 2(c)). With a further increase of $E_{0}/\Omega$, the effect of the dynamical localization point becomes evident, where $d_{x^{2}-y^{2}}+id_{xy}$ order is stabilized again, but now with the $d_{xy}$ component being dominant ($d\mathrm{SC}2$). This phase appears in the vicinity of the magnetic instability towards the stripe AFM order ($\mathrm{AFM}2$), rather than $\mathrm{AFM}1$. All orders are found to be topologically nontrivial (see SM [55]).
Discussion.
@@ Appendix A Floquet Green’s function formalism
$\Sigma^{\mathrm{R/A}}_{\mathrm{res};nn^{\prime}}(\omega)$ $=\mp i\gamma\delta_{nn^{\prime}}\;,$ (21)
@@ Appendix D Floquet engineering the effective pairing vertex
Figure 5: The real space components of the pairing vertex as a function of drive amplitude for different drive frequencies. The lattice constant is set to $a=1$. The first row shows the $n=0$ component, i.e. the time-averaged vertex. The second row shows the $n=1$ component that describes oscillations of the vertex. (a) and (d) show the onsite interaction which is repulsive. (b) and (e) show the nearest neighbor interaction, which is attractive in the time average, facilitating $d_{x^{2}-y^{2}}$ order. (c) and (f) show the next-nearest neighbor interaction that remains repulsive in the time average but becomes attractive in the $n=1$ component. The magnitude of the $n=1$ terms is strongly suppressed by $\Omega$.
To further understand the mechanism that generates the $id_{xy}$ symmetry components, it is insightful to investigate the real space structure of the effective pairing vertex, see Fig. 5. The bonds on which the effective interaction becomes attractive (negative) determine the gap symmetry. For $d_{xy}$-symmetry, an attractive interaction on the next-nearest neighbor bonds is required. The interaction on the next-nearest neighbor bonds only becomes attractive in the higher order $n=1$ components of the vertex. Therefore, oscillations in the effective interaction that are induced through the drive give the gap its $d+id$ character. This also explains the suppression of the $id_{xy}$ components, as these higher order terms are suppressed by $\Omega$.
@@ Appendix F Floquet amplitude expansion
We evaluate Equation˜66 for $\omega=0$ and $\mathbf{q}=\mathbf{q}_{C}$, where $\mathbf{q}_{C}$ is determined from the exact calculation described in the main text. According to our analysis, $\mathbf{q}_{C}$ can also be accurately obtained from Equation˜66 up to $E_{0}/\Omega\leq 0.4$ for $\Omega=2$, $E_{0}/\Omega\leq 0.8$ for $\Omega=4$, $E_{0}/\Omega\leq 2$ for $\Omega=6$, and $E_{0}/\Omega\leq 2.2$ for $\Omega=8$.